A watermelon is suspended from a spring and causes the spring to expand 0.54 m. If the spring constant is 131 N/m, what is the mass of the water melon?
1
Expert's answer
2020-12-04T12:17:46-0500

According to the Hook's law, the force acting on the spring is equal to:


"F = kx"

where "k = 131N\/m" is the spring constant, and "x = 0.54m" is the expansion of the spring. On the other hand, the force acting on the spring from the water melon is basically its weight. Thus


"F = mg"

where "m" is the mass of the water melon and "g = 9.81m\/s^2" is the gravitational acceleration. Substituting this expression into the Hook's law, obtain:


"mg = kx\\
m = \dfrac{kx}{g} = \dfrac{131\cdot 0.54}{9.81} \approx 7.21kg"

Answer. 7.21 kg.